OpenStudy (anonymous):

. To produce a textbook, suppose the publisher spent $125,000 for typesetting and $7.50 per book for printing and binding. The total cost (C) to produce and print n books can be written as an equation C(n) = 125,000 + 7.5n. If the average cost is the total cost divided by the number of books printed, find the average cost of producing 10,000 books.

hero (hero):

c(n)/n = c(10,000)/10000

OpenStudy (anonymous):

If the average cost is the total cost divided by the number of books printed, you get \[A(n)=\frac{C(n)}{n}=\frac{125000+7.5n}{n}\]

OpenStudy (anonymous):

?

hero (hero):

\[A(10,000) = \frac{c(10000)}{10000}=\frac{125000+7.5(10000)}{10000}\]